# Structured Unrestricted-Rank Matrices for Parameter Efficient Fine-tuning

### v2

@@ 4.2 Low-rank versus Circulant and Toeplitz matrices
In this section, we focus on the difference in approximation qualities between low rank matrices and the circulant and Toeplitz matrices by using a fixed parameter budget. We use the following two settings:
Approximating Symmetric Positive Definite Matrices. $\mathbf{M}\in\mathbb{R}^{n\times n}$ ($n=50$) with $L^{2}$-normalized rows is used for this experiment. We compare the errors to approximate $\mathbf{M}$ using circulant, (symmetric) Toeplitz, low-rank matrices, and Kronecker product of $2$ matrices. A fixed parameter budget of $n=50$ is chosen and this experiment is repeated $10$ times. Our results are presented in Fig. 1, Left. For each of these $10$ trials, we observe that the low rank matrix has the highest approximation error followed by the Kronecker product of two matrices, the circulant, and the Toeplitz matrices (see more details in Appendix A.5).
@@
Figure 5: The approximation capabilities of different LDRMs. The $y$-axis depicts the relative Frobenius norm error ${\|\mathbf{A}-\mathbf{M}\|_{\mathrm{F}}}{/\|\mathbf{M}\|_{\mathrm{F}}}$ between the groundtruth matrix $\mathbf{M}\in\mathbb{R}^{100\times 100}$ and the approximator $\mathbf{A}$. Left Column Top: Approximating random Gaussian matrix $\mathbf{M}$ with matrices $\mathbf{W}(\mathbf{G},\mathbf{H})$ with different $r$ (LDR: $r$). Left Column Middle: Approximating near-low-rank matrices $\mathbf{M}$ replacing random Gaussian matrices and smaller values of $r$. Left Column bottom: Similar setup to approximate near-low-intrinsic-rank matrices $\mathbf{M}$. Right Column: We perform analogous studies with circulant and Toeplitz matrices, where the ground truth has low rank or low-intrinsic rank.
@@
Fitting a simple toy dataset : We create a pinwheel dataset with 5 spokes (adding a little Gaussian noise) (Fig : 8 Left). Then we try to fit a simple neural network with one hidden layer with the matrix of size $64\times 64$. We then replace this matrix with a rank $1$ LoRA layer, a circulant layer, a symmetric Toeplitz and a Toeplitz layer making sure that all of these models have the same number of training parameters. We observe that the LoRA layer struggles to fit the data whereas the low displacement rank matrices show similar performance to that of the baseline (Fig 4). These results show the impressive expressive power of these matrices. Additional details are presented in Appendix B and additional experiments are presented in the Appendix B.2 (see Fig 8).
We thus conclude that LDRMs with particularly low displacement rank serve as good approximators for various matrices. This power translates well to various downstream tasks as will be confirmed by our experimental results, with circulant matrices performing especially well (see: Sec. 6).
@@ 5 Integration of SURMs with PEFT
Motivated by the results of the previous section, we are now ready to use SURMs as drop-in replacements for various PEFT methods. In this section, we discuss the integration of SURMs in LoRA and relegate the discussion about SURMs in Adapters to the Appendix. A.3. Our design choices take into account the trade-off between expressivity and training efficiency.
5.1 Transforming LoRA with SURMs
LoRA [26] proposes an adaptation of the weight matrix via low rank updates. More concretely, if $\mathbf{W}$ is the pretrained weight, then the updated $\widehat{\mathbf{W}}=\mathbf{W}+\alpha\Delta\mathbf{W}$, where $\Delta\mathbf{W}=\mathbf{AB}^{\top}$ and $\mathbf{A}\in\mathbb{R}^{m\times r},\mathbf{B}\in\mathbb{R}^{n\times r}$ for $r\ll\text{min}(m,n)$ and $\alpha$ is a fixed scaling parameter. For efficient training, $\Delta\mathbf{W}$ needs to be initialized as a zero-matrix. Importantly, in LoRA it is done by choosing $\mathbf{B}$ to be a random matrix and initializing $\mathbf{A}$ to be the zero matrix. In this work, we propose learnable SURMs as $\mathbf{AB}$-replacements for $\Delta\mathbf{W}$. For simplicity of notation, we will assume $m=n$.
Circulant Matrices : Let $\mathbf{C}\in\mathbb{R}^{n\times n}$ be a circulant matrix which we parameterize by an $n$-dimensional vector $\mathbf{r}$ encoding its first row. Instead of assigning: $\Delta\mathbf{W}=\mathbf{C}$, we instead take: $\Delta\mathbf{W}=\mathbf{C}_{1}\odot\mathbf{C}_{2}$ with the corresponding vectors $\mathbf{r}_{1}\odot\mathbf{r}_{2}$, where $\odot$ is the Hadamard product (element-wise multiplication). Following the strategy applied in LoRA, we initialize $\mathbf{C}_{1}$ as a zero-vector and $\mathbf{C}_{2}$ as a random-vector. Opting for Hadamard products over conventional matrix products is primarily driven by the advantage of utilizing efficient multiplication only once. The construction of Hadamard products which is O($n$) is quicker than the process involved in efficient multiplication (which is O$(n\text{log}(n))$. Additionally, this approach does not compromise the expressiveness of the network. In both scenarios, the result is a circulant matrix, as the product of two circulant matrices yields another circulant matrix.
Toeplitz Matrices : Let $\mathbf{T}\in\mathbb{R}^{n\times n}$ be a Toeplitz matrix which we parameterize by an $n$-dimensional vector $\mathbf{r}$ encoding its first row and an $n$-dimensional vector $\mathbf{c}$ encoding its first column (thus in total $12n-$ 1 parameters). To address the initialization challenge, we apply two Toeplitz matrices where one is initialized randomly while the other is the zero matrix (same strategy as before).
@@ 5.1 Transforming LoRA with SURMs
Note that the product of two Toeplitz matrices may not be Toeplitz. We define $\Delta\mathbf{W}=g(\mathbf{T_{2}},g(\mathbf{T_{1}},\mathbf{x}))$, where $\mathbf{T_{1}},\mathbf{T_{2}}$ are two Toeplitz matrices, and $g$ is the operator that allows efficient matrix-vector multiplication with Toeplitz matrices (see Appendix A.2 ). This formulation leads to the $24n-$ 2 trainable parameters. To further reduce this number, we constrain the $\mathbf{T_{1}},\mathbf{T_{2}}$ to be symmetric, reducing the total number of trainable parameters to $2n$.
Kronecker Product of Matrices : In this case, we write $\Delta\mathbf{W}=\mathbf{A}\otimes\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{r_{1}\times r_{2}},\mathbf{B}\in\mathbb{R}^{\frac{n}{r_{1}}\times\frac{n}{r_{2}}}$. The hyperparameters $r_{1},r_{2}$ allow us to control not only the number of trainable parameters but also the rank of $\Delta\mathbf{W}$. In contrast to low-rank matrix updates, one can create matrices $\Delta\mathbf{W}$ of fairly large ranks while keeping the number of trainable parameters small (see Appendix A.1) We follow the same initialization strategy as before.
Table 2: Results on the VTAB-1k benchmark. Baseline numbers are taken from [29] and [49]. Best numbers are highlighted in bold and the second-best numbers are underlined.
@@
Figure 6: Low Resource Training. Accuracy of SURM Clip-ViT models as a function of the fraction of the training data. The results show that SURM can achieve comparable accuracy with as low as $\sim 2\%$ of the training data for easier tasks like CIFAR10 and $\sim 20\%$ for harder tasks like SUN397.
In all the mentioned scenarios, it is possible to increase the number of training parameters by relaxing the matrix structure. This can involve incorporating more circulant or Toeplitz matrices in the product chains, utilizing asymmetric Toeplitz matrices, adjusting the sizes of factors in the Kronecker product, or employing sums of such matrices. Another way to enhance layer expressiveness is to experiment with combinations of different LDRMs, such as a mix of circulant and skew-circulant matrices. We leave this exploration for future investigation.
Figure 7: Left: Average tradeoff between accuracy and parameter numbers of various PEFT methods. Results are measured across 5 image datasets using ViT-B. SURMs appear in the top right corner and perform best among various strong baseline methods. The complete result is shown in Table 1. Right: Average performance reported across GLUE benchmark task (8 tasks). SURMs appears at the top right corner indicating SURM achieves better performance with a fraction of trainable parameters. The detailed results for individual tasks are reported in Appendix B.2 (Table 5).
@@ 6.1 Vision Experiments
We test SURM on several vision datasets: CIFAR10, CIFAR100 [38], SUN397 [76], DTD [16] and STL10 [17]. We focus on the LoRA setting and adapt $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}$ matrices. ViT-B/16 [35] and Clip-ViT-B/16 [61] are used as base models. Results using ViT-base are presented in Table 1 (left). SURM consistently outperform $\mathbf{12}$ baseline methods that use up to 10 times as many parameters. On three out of the five tasks under consideration, SURMs emerge as the top performers, surpassing LoRA by a margin as substantial as 5-$\textbf{7}\%$. Meanwhile, SURMs maintain competitiveness on the remaining two tasks. Results using Clip-ViT are presented in Table 1 (right). In this setting, our methods are one of the top two methods among all $\mathbf{12}$ baselines on all 5 tasks. Furthermore, our method is characterized by an exceptionally small number of trainable parameters, resulting in 3.65x reduction compared to LoRA, and a 2.4x reduction compared to LoRA-Fix.
@@
Next, we evaluate SURM in low resource setting using VTAB-1k datasets [81] using the ViT as the base model against $12$ strong baselines (Table 2). VTAB-1k is a diverse collection of vision datasets with only $1000$ examples for training, and in this work we focus on the NATURAL and SPECIALIZED subsets of VTAB. Among the $11$ tasks examined, our approaches are one of the top 2 methods on $\mathbf{10}$ datasets, competitive on $2$ others. Additional results on ImageNet can be found in Appendix B.2.
@@
Finally, we evaluate the efficacy of SURM in low data regime. We use the Clip-ViT model and train on a fraction of training data, showing the results in Figure 6. We find that among our three proposed SURM methods, Circulant works best in low data regime. Moreover, we match the full fine-tuning accuracy on the entire training set with only a small fraction of the data. For more challenging datasets such as Sun397, we achieve comparable accuracy using approximately $\sim 20\%$ of the training data, while for datasets like CIFAR10 and STL10, only about $2\%$ of the training data is required.
As mentioned earlier, contrary to our work, all the previous Kronecker-based methods constrain the Kronecker adaptations to be low rank. For a detailed discussion about the difference between our Kronecker adaptation and existing methods, see Appendix B.3.
NLP Experiments. We exhaustively evaluate our SURM models on the GLUE benchmark [71]. We compare with different adapter baselines and 11 other PEFT techniques. These include Bert-base-uncased [20] full finetuning, Adapter (Houlsby) and Adapter (Pfeiffer) whose numbers are taken from [57]. BiTFit results are taken from [80] (except QQP numbers which are obtained from [9]) and the numbers for $AA$-adapters from [52]. Prefix, Serial, AdaMix, UniPELT, Parallel, MAM, and AutoPEFT numbers are taken from [86]. The results for the remaining baselines are replicated by us. More experimental details can be found in Appendix B.
Image Segmentation. Next, we focus on the extremely challenging task of medical image segmentation by using the Synapse multi-organ segmentation dataset  [^1] https://www.synapse.org/#!Synapse:syn3193805/wiki/217789. Details of the dataset is presented in Appendix B. The Segment-Anything-Model (SAM) [33] is used as the foundation model for this task. We follow the training details in [82]. and adapt the $\mathbf{Q},\mathbf{V}$ in ViT-B image encoder in the SAM. Finally, in this small data regime, we use the Circulant variant as it is our most performant variant (Fig. 6). We report the Dice similarity coefficient (DSC) metric for each of the $8$ organ segmentations as well as the average DSC score for all (higher is better). For a fair comparison, we include LoRA with rank $\mathbf{1}$, matching the exact parameter count of Circulant. The results are presented in Table 3. We report the baseline performance from [82]. SURMs compare favorably with specialized architectures developed for medical imaging like U-Net, Attention U-Net, Transformer-based U-Net, and the Swin U-Net even though they have significantly higher number of training parameters than our method.
@@
The full results for all the SURM methods are shown in Appendix Table 5. For brevity, we summarize the average performance across 8 tasks for SURM-adapters, as compared to all 11 baselines, in Fig 7 Right. SURM achieve much better performance at a fraction of the parameters.
Our SURM-LoRA method outperforms the baseline LoRA. For a fair comparison, we select the ranks in LoRA such that the total number of trainable parameters is consistent across all methods (Table 5). We perform additional ablations in Appendix 14(a). Finally, we analyze the representations learnt by SURMs (Appendix B.4). Our main finding is that the LoRA learnt weights are very similar to the pre-trained weights whereas SURMs explore a larger space (an observation similar to [87]).

### v3

@@ 4.1 Comparing $\mathbf{W}(\mathbf{G},\mathbf{H})$ with Circulant and Toeplitz Matrices
In Figure 5, we report the relative Frobenius norm error during training for different settings. In Fig 5 (top left), we use $\mathbf{W}(\mathbf{G},\mathbf{H})$ with different $r$ (rank of $\mathbf{F}$ in Eq. 1) is used to approximate random matrices. While the best approximations are achieved for larger values of $r$ (specifically, $r=20$), it is interesting to note that the final error does not decrease monotonically with increasing $r$. For the remaining class of matrices $\mathbf{M}$, which are close to low-rank and therefore easier to approximate, we experiment with smaller values of $r$ and report the results in Figure 5 (left column, middle and bottom). In this case, we observe that the three top-performing approximators $\mathbf{W}(\mathbf{G},\mathbf{H})$ were trained with $r=1,2,4$. These results indicate that for more structured ground truth matrices (even if they are not necessarily low-rank), LDRMs with a very low rank for the corresponding $\mathbf{F}$ are sufficient.
Motivated by the results showing that LDRMs with low $r$ can serve as effective approximators, we use circulant and Toeplitz matrices to approximate near-to-low-rank and low-intrinsic-rank matrices. In the three plots shown in Fig. 5 (right column), we observe that approximations using Toeplitz matrices (using twice as many parameters as circulant matrices) offer negligible gains and are only beneficial in the near-low-rank case. For the low-intrinsic case, circulant matrices outperform Toeplitz ones. Overall, circulant matrices with few parameters achieve strong performance in this setting.
@@
Figure 5: Illustration of the approximation capabilities of different LDRMs. The $y$-axis depicts the relative Frobenius norm error ${\|\mathbf{A}-\mathbf{M}\|_{\mathrm{F}}}{/\|\mathbf{M}\|_{\mathrm{F}}}$ between the groundtruth $\mathbf{M}$ and the approximator $\mathbf{A}$. (Left Column Top): We approximate a random Gaussian matrix $\mathbf{M}$ with matrices $\mathbf{W}(\mathbf{G},\mathbf{H})$ using different $r$ (LDR: $r$). (Left Column Middle) We approximate near-low-rank matrices $\mathbf{M}$ using smaller values of $r$. (Left Column Bottom): Similar setup to approximate near-low-intrinsic-rank matrices $\mathbf{M}$. (Right Column): We perform analogous studies with circulant and Toeplitz matrices, where the ground truth has low rank or low-intrinsic rank.
4.2 Comparing Low Rank with Circulant and Toeplitz Matrices
In this section, we focus on the difference in approximation qualities between low rank matrices and the circulant and Toeplitz matrices under a fixed parameter budget. We use the following settings:
Approximating Symmetric Positive Definite Matrices. We use a PSD matrix $\mathbf{M}\in\mathbb{R}^{50\times 50}$ with $L^{2}$-normalized rows in this experiment. We compare the errors to approximate $\mathbf{M}$ using circulant, (symmetric) Toeplitz, low-rank matrices, and Kronecker product of two matrices. We use a fixed parameter budget and repeat this experiment $10$ times. We report the results in Figure 1 (left). For each of these $10$ trials, we observe that the circulant and Toeplitz achieve the lowest error, therefore the best approximation quality (see more details in Appendix A.7).
@@ 4.2 Comparing Low Rank with Circulant and Toeplitz Matrices
Fitting a Toy Dataset. We create a synthetic pinwheel dataset with 5 spokes as shown in Figure 9 (left). We fit this dataset using a simple neural network with one hidden layer with a matrix $\mathbf{W}\in\mathbb{R}^{64\times 64}$. In this experiment, we replace $\mathbf{W}$ with a rank $1$ LoRA, a circulant, a symmetric Toeplitz, and a Toeplitz matrix ( with all parameterizations having the same number of training parameters). In Fig 4), we report the training loss curves of this experiment. We observe that the LoRA layer struggles to fit the data whereas the LDRMs show similar performance to full fine-tuning (Fig 4). These results show the impressive expressive power of these matrices. Therefore, we conclude that LDRMs with particularly low displacement rank serve as good approximators for various matrices. We perform additional experiments and provide more details in Appendix B.
@@ 5 Integration of SURMs with PEFT
Motivated by the results from the previous section, we use SURMs as drop-in replacements for various PEFT methods. In this section, we present the integration of SURMs in two popular classes of PEFT methods: LoRA and Adapters.
5.1 Integration of SURMs in LoRA
LoRA [26] uses a low rank matrix to parameterize the weight matrix updates. Formally, given a pre-trained weight matrix $\mathbf{W}$, then the updated matrix is $\widehat{\mathbf{W}}=\mathbf{W}+\alpha\Delta\mathbf{W}$, where $\Delta\mathbf{W}=\mathbf{AB}^{\top}$ and $\mathbf{A}\in\mathbb{R}^{m\times r},\mathbf{B}\in\mathbb{R}^{n\times r}$ for $r\ll\text{min}(m,n)$ and $\alpha$ is a fixed scaling parameter. For efficient training, $\Delta\mathbf{W}$ needs to be initialized as a zero-matrix. LoRA performs this by choosing initializing $\mathbf{A}$ to be the zero matrix and $\mathbf{B}$ to be a random matrix. In this work, we propose to parameterize $\Delta\mathbf{W}$ using structured unrestricted rank matrices. Next, we provide the details of parameterizing $\Delta\mathbf{W}$ using different SURM matrices (assuming $m=n$ for simplicity).
Circulant Matrices. In this setting, we parameterize the updates as: $\mathbf{W}=\mathbf{C}_{1}\odot\mathbf{C}_{2}$, where $\mathbf{C}_{i}\in\mathbf{R}^{n\times n}$ are circulant matrices encoded using a $n$-dimensional vector, $\mathbf{r}_{i}\in\mathbb{R}^{n}$.
We use Hadamard products ($\mathbf{C}_{1}\odot\mathbf{C}_{2}$) instead of conventional matrix products as it can be computed efficiently. The construction of Hadamard products which is O($n$) is quicker than the process involved in efficient multiplication (which is O$(n\text{log}(n))$. To enable zero-initialization, we initialize $\mathbf{C}_{1}$ as a zero-vector and $\mathbf{C}_{2}$ as a random-vector. Additionally, this approach does not compromise the expressiveness of the network, as the result of the Hadamard product is also a circulant matrix.
Toeplitz Matrices. Similar to the previous setting, we use two Toeplitz matrices to parameterize: $\Delta\mathbf{W}=g(\mathbf{T_{2}},g(\mathbf{T_{1}},\mathbf{x}))$, where $\mathbf{T_{1}},\mathbf{T_{2}}$ are Toeplitz matrices, and $g$ is the operator that allows efficient matrix-vector multiplication with Toeplitz matrices (see Appendix A.3 ). Each Toeplitz matrix $\mathbf{T}\in\mathbb{R}^{n\times n}$ is parameterized using an $n$-dimensional vector $\mathbf{r}$ encoding its first row and an $n$-dimensional vector $\mathbf{c}$ encoding its first column ($12n-$ 1 total parameters). This formulation leads to the $24n-$ 2 trainable parameters. To further reduce this number, we constrain the $\mathbf{T_{1}},\mathbf{T_{2}}$ to be symmetric, reducing the total number of trainable parameters to $2n$. To enable zero-initialization, we initialize $\mathbf{T}_{1}$ as a zero matrix and $\mathbf{T}_{2}$ is randomly initialized.
@@ 5.1 Integration of SURMs in LoRA
Kronecker Product of Matrices. In this setting, we parameterize: $\Delta\mathbf{W}=\mathbf{A}\otimes\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{r_{1}\times r_{2}},\mathbf{B}\in\mathbb{R}^{\frac{n}{r_{1}}\times\frac{n}{r_{2}}}$. The hyperparameters $r_{1},r_{2}$ allow us to control the trainable parameter count and the rank of $\Delta\mathbf{W}$. In contrast to low-rank matrix updates, we can create matrices $\Delta\mathbf{W}$ of fairly large ranks while keeping the number of trainable parameters small (see Appendix A.2). To enable zero-initialization, we set $\mathbf{A}$ as a zero matrix and $\mathbf{B}$ as a random matrix.
Table 2: Results on the VTAB-1k benchmark. Baseline numbers are taken from [30] and [50]. Best numbers are highlighted in bold and the second-best numbers are underlined. We observe that SURM is one of the top-performing PEFT methods on almost all datasets.
@@
Figure 6: Low Resource Training. Accuracy of SURM CLIP-ViT models as a function of the training data fraction. The results show that SURM can achieve comparable accuracy with as low as $\sim 2\%$ of the training data for easier tasks like CIFAR10 and $\sim 20\%$ for harder tasks like SUN397.
In all the above settings, it is possible to increase the number of training parameters by relaxing the structure of the matrix, $\Delta\mathbf{W}$. This can be performed by introducing more matrices in the product chains, utilizing asymmetric Toeplitz matrices, adjusting the sizes of factors in the Kronecker product, or employing sums of such matrices. Another way to enhance layer expressiveness is by experimenting with combinations of different LDRMs, such as mixing circulant and skew-circulant matrices. A broad class of matrices, including low-rank ones, can be represented as sums of these matrices (see Theorem A.2 and the subsequent discussion).
5.2 Integration of SURMs in Adapters
Adapters [25] are small bottleneck networks into Transformer layers as shown below:
${}\mathbf{Y}=\mathbf{X}+\sigma(\mathbf{XB})\mathbf{A},$ (4)
where $\sigma(\cdot)$ is a non-linear activation function, $\mathbf{X}\in\mathbb{R}^{b\times s\times n}$ represents input to the layer ($b$: batch size, $s$: sequence length), $\mathbf{A}\in\mathbb{R}^{r\times n},\mathbf{B}\in\mathbb{R}^{n\times r}$ are low-rank matrices ($r\ll n$) and $\mathbf{Y}$ is the output of the layer. For simplicity, layer norms and bias terms are not included in the equation. We use SURMs as a drop-in replacement for matrices $\mathbf{A}$ and $\mathbf{B}$. Next, we will provide details of integrating SURMs within the adapter setting. Additional details are provided in Appendix A.5.
Circulant Matrices. In this setting, we apply two circulant matrices $\mathbf{C}_{1},\mathbf{C}_{2}$ (encoded by $\mathbf{r}_{1},\mathbf{r}_{2}$), resulting in the adapter block: $\mathbf{Y}=\mathbf{X}+\sigma(f(\mathbf{r_{1}}\odot\mathbf{r_{2}},\mathbf{X}))+\mathbf{b}$, where $f$ is an operator that efficiently computes the matrix multiplication between input $\mathbf{X}$ and the circulant matrix encoded by the vector $\mathbf{r}_{1}\circ\mathbf{r}_{2}$ (Appendix A.3). To enable zero-initialization, the vector $\mathbf{r}_{1}$ is initialized randomly while $\mathbf{r}_{2}$ and $\mathbf{b}$ are initialized as zero vectors.
Toeplitz Matrices. In this setting, we use two symmetric Toeplitz matrices $\mathbf{T_{1}},\mathbf{T_{2}}$, where $\mathbf{T_{1}}$ and $\mathbf{b}$ is initialized as a zero vector and $\mathbf{T_{2}}$ is initialized randomly. We then define the adapter layer as: $\mathbf{Y}=\mathbf{X}+\sigma(g(\mathbf{T_{1}},g(\mathbf{T_{2}},\mathbf{X})))+\mathbf{b}$, where $g$ is an operator that efficiently computes the matrix multiplication between an input $\mathbf{X}$ and a Toeplitz matrix.
Kronecker Product of Matrices. In this case, we rewrite Equation 4 as: $\mathbf{Y}=\mathbf{X}+\sigma(\mathbf{X}(\mathbf{B}\otimes\mathbf{A}))+\mathbf{b}$, where $\mathbf{B}\otimes\mathbf{A}$ is the Kronecker product. In this case, $\mathbf{B}$ is initialized randomly and $\mathbf{A}$ and $\mathbf{b}$ are initialized by zeros. In all experiments using SURM-adapters, $\sigma(\cdot)$ is the GeLU non-linearity.
Figure 7: Left: Tradeoff between performance and parameter count for various PEFT methods. We report the average results across 5 image datasets using ViT-B (complete results in Table 1). Right: Average performance across GLUE benchmark (see complete results in Table 5). SURMs appear in the top right corner and perform best among various strong baseline methods in both settings.
@@ 6 Experiments
Image Classification Experiments. We evaluate SURM on several vision datasets: CIFAR10, CIFAR100 [39], SUN397 [79], DTD [16] and STL10 [17]. We experiment using ViT-B/16 [36] & Clip-ViT-B/16 [63] as base models and inject trainable parameters $\mathbf{Q},\mathbf{K},\mathbf{V}$ matrices in the LoRA setting. We report the results using ViTbase are presented in Table 1 (left). We observe that SURM consistently outperforms $\mathbf{12}$ baseline methods (that use up to 10x parameters). On three out of the five tasks, SURMs emerge as the top performers, surpassing LoRA by a margin of up to 5-$\textbf{7}\%$ while achieving competitive performance on the remaining tasks. We report the results using Clip-ViT in Table 1 (right). In this setting, SURM is among the top two methods across all 5 tasks. SURM also uses fewer trainable parameters, reducing them by 3.65x compared to LoRA and 2.4x compared to LoRA-Fix.
@@
Low Data Regime. We evaluate SURM in low data regime using VTAB-1k datasets [85] and the ViT model. VTAB-1k is a diverse collection of vision datasets with only $1000$ training examples. We focus on the Natural and Specialized subsets of VTAB. In Table 2, we observe that SURMs are among the top 2 methods on $\mathbf{10}$ datasets while being competitive on the remaining tasks.
@@
We evaluate different variants of SURM and train it on a varying fraction of data on 5 datasets using Clip-ViT model. We report the results in Figure 6. We observe that the circulant SURM works best in low data regime. Furthermore, SURM achieves the performance of full fine-tuning trained on the entire dataset with only a small fraction of the data. For more challenging datasets like Sun397, we achieve comparable accuracy using approximately 20% of the training data, while for datasets such as CIFAR10 and STL10, only about 2% is needed.
Large Data Regime. We perform experiments to show that SURM generalizes well in large data regimes. On ImageNet [18] and iNat2021 [72], SURM achieves performance comparable to full fine-tuning while using only $\mathbf{0.06}\%$ of the training parameters (see detailed results in Appendix B.2).
@@
NLP Experiments. We extensively evaluate SURM models on the GLUE benchmark [74] using BERTbase [20]. We compare with different adapter baselines and 11 other PEFT techniques. These include full-finetuning, Adapter (Houlsby) and Adapter (Pfeiffer), among others (we report the corresponding results from [59]). BiTFit results are taken from [84] (except QQP numbers which are obtained from [9]) and the numbers for $AA$-adapters from [53]. Prefix, Serial, AdaMix, UniPELT, Parallel, MAM, and AutoPEFT numbers are taken from [90]. The results for the remaining baselines are replicated by us. More experimental details can be found in Appendix B.
For brevity, we summarize the average performance across 8 tasks for SURM-adapters and compare it to 11 baselines, in Fig 7 (right). We observe that SURM achieve much better performance while using a fraction of the parameters (the complete results are reported in Appendix Table 5). We also observe that SURM (integrated into LoRA) outperforms the baseline LoRA, under the same parameter budget. This shows the effectiveness of using structured matrices as a drop-in replacement for low rank matrices used in LoRA. We further analyze the representations learnt by SURMs and LoRAs (Appendix B.4). We find that LoRA learns weights that are very similar to the pre-trained weights whereas SURM is able to explore a larger parameter space (an observation similar to [91]).
Figure 8: Fitting the UUID dataset using Llama-2-7b. We fit the data using various SURM-based PEFT methods and LoRA.
Large-scale Experiments. In this setting, we integrate SURMs in LLMs. Specifically, we use matrices of the form $\mathbf{W}(\mathbf{G},\mathbf{H})$ (as described in Eqn 2) to increase the number of training parameters. We use the experimental setup introduced in [29], where the LLM tries to fit a dataset of UUID pairs using the Llama-2-7B model [69]. This was shown to be a challenging task (UUID prediction is significantly different from the pre-training tasks) that requires higher rank values in LoRA. We report the results in Fig. 8, demonstrating that SURMs is able to fit the data, whereas other methods struggle to do so (see more details in Appendix 9).
Image Segmentation. Next, we focus on the extremely challenging task of medical image segmentation using Synapse multi-organ segmentation dataset [82]. Segment-Anything-Model (SAM) [34] is used as the foundation model for this task. We follow [86] and adapt the $\mathbf{Q},\mathbf{V}$ in ViTbase image encoder in SAM. Finally, in this low data regime, we use Circulant variant of SURM as it is the best performing variant (Fig. 6). We report the Dice similarity coefficient (DSC) metric for each of the $8$ organ segmentations and their average (higher is better). For a fair comparison, we include LoRA with rank $\mathbf{1}$, matching the exact parameter count of Circulant. The results are presented in Table 3. We report the baseline performance from [86]. SURMs compare favorably with specialized architectures developed for medical imaging like U-Net, Attention U-Net, Transformer-based U-Net, and the Swin U-Net even though they have significantly higher number of training parameters than our method (see details in Appendix 9).